OpenStudy (anonymous):

EVALUATE THE EXPONENTIAL FUNCTION FOR THREE POSITIVE VALUES OF X,THREE NEGATIVE VALUES OF X,AND AT X=0.SHOW YOUR WORK.USE THE RESULTING ORDERED PAIRS TO PLOT THE GRAPH.STATE THE DOMAIN AND RANGE OF THE FUNCTIONATTACH GRAPH AS LISTED AND REQUIRED.F(X)=E^-X-1

OpenStudy (anonymous):

please help me out in this......

OpenStudy (anonymous):

\[f(x) = e ^{-x}-1\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

yes this is the question?

OpenStudy (anonymous):

okay, substitute and three positive values of x, say 1,2,3 and three negative values, say -1,-2,-3 in f(x)

OpenStudy (anonymous):

so for x =1, f(x) = f(1) = (e^-1) -1 use your calculator or google to find the value of e^-1 and subtract 1 from that to get f(1)

OpenStudy (anonymous):

just go to google.com and type e^-1. you will get a value. try it out

OpenStudy (anonymous):

according to google calculator e^-1 is 0.367879441. so (e^-1) -1 = -0.632120559 so at x = 1, f(x) = -0.632120559

OpenStudy (anonymous):

similarly find out values of f(x) for x =2,3 and x = -1,-2,-3

OpenStudy (anonymous):

use those values and make a graph such as this: f(x) | | | | | | | |----------------------->x

OpenStudy (anonymous):

iam new to this..thts why getting problems

OpenStudy (anonymous):

okay I understand. But now I have shown you the procedure. Just follow it. Its easy. You can do it. Did you try typing it in google?

OpenStudy (anonymous):

yes i did but i dint got

OpenStudy (anonymous):

okay. do you have a scientific calculator with you?

OpenStudy (anonymous):

did you type in e^-1 exactly as I typed it? just copy e^-1 and paste it in google. and press enter.

OpenStudy (anonymous):

yeah i did the same

OpenStudy (anonymous):

i got now

OpenStudy (anonymous):

okay what did you get?

OpenStudy (anonymous):

can you help me in another question?

OpenStudy (anonymous):

i got the same which you got

OpenStudy (anonymous):

okay good, just do the same for other numbers and you will get your answers.

OpenStudy (anonymous):

Yes, I can help you. But your new problem separately.

OpenStudy (anonymous):

post your new problem separately.

OpenStudy (anonymous):

ok can i post it here

OpenStudy (anonymous):

no.. post it separately. this post is too long already

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I dont see it zehera.

OpenStudy (anonymous):

its not coming there

OpenStudy (anonymous):

i think that i have poted it before thts y it not coming once again

OpenStudy (anonymous):

eveluate the exponentaial eqution forthree positive valu f x,and three negative values of x,and at x=0.transform the second expression into the equivalent logarithm equation and evalaute the logarithm equation for thre values of x that are greater than 1,three values of x that are between 0 and1,and at x=1.show your work.use the resulting ordered pairs to plot the graph of each function.Y=5x^-2,X=5y^-2

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so Y = 5x^-2 is equation 1 and X = 5Y^-2 is equation 2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay what is x^-2 ?

OpenStudy (anonymous):

on google i got this but ith another equation

OpenStudy (anonymous):

im getting positive values but for negative vl\alues iam getting confused

OpenStudy (anonymous):

okay, ;ets make it simpler then. what is x^-2 ?

OpenStudy (anonymous):

ok it would be greatful

OpenStudy (anonymous):

okay. but what is x^-2?

OpenStudy (anonymous):

its given that Y=5X^-2

OpenStudy (anonymous):

no, tell me what x^-2 means. I dont want maths. Just tell me in plain english. What do you think x^-2 means?

OpenStudy (anonymous):

I DONT KNOW?

OpenStudy (anonymous):

okay. \[x ^{-2} = 1/x ^{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so, \[5x ^{-2} = 5/x ^{2}\]

OpenStudy (anonymous):

that means, \[Y = 5/x ^{2}\]

OpenStudy (anonymous):

so now substitute x = 1,2,3 and -1,-2,-3 and get Y values. what is the value of Y at 0?

OpenStudy (anonymous):

1 i think

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

any value at log0 is 1

OpenStudy (anonymous):

you are just asked to evaluate the first equation. Why are you taking log?

OpenStudy (anonymous):

ok you explain me how to do

OpenStudy (anonymous):

if we take one value as 2 thn y=5/4

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok like this we have to take the other two values

OpenStudy (anonymous):

right. But the question also asks for x =0 . so what happens if you take another value as 0?

OpenStudy (anonymous):

dont lknow

OpenStudy (anonymous):

just put 0 in the equation. see what you get. y = 5/0. what is anything divided by 0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so we have to take three values as 0,1,2

OpenStudy (anonymous):

no, anything multiplied by 0 is 0. anything divided by zero is infinity.

OpenStudy (anonymous):

No, pay attention zehera. the problem asks for three positive values. so you use any three positive values. Similarly use three negative values. then use x = 0 which is neither positive nor negative.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so totally you have 7 values, three positive, three negative and one 0.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay. now second equation.

OpenStudy (anonymous):

while taking negative values we get 5/-4

OpenStudy (anonymous):

what does the problem say you should do with the second equation?

OpenStudy (anonymous):

if x = -2, what is x ^2?

OpenStudy (anonymous):

x^2 = x * x. so, if x = -2, x * x = ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

so then for x = -2, Y = ?

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

x = -2. Y = 5/(-2)^2

OpenStudy (anonymous):

but you just said (-2)^2 = 4 so Y = 5/4.

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

but here the poer is -2 na

OpenStudy (anonymous):

-2^-2=?

OpenStudy (anonymous):

are you talking about equation 1 or equation 2?

OpenStudy (anonymous):

equation2

OpenStudy (anonymous):

okay i was talking about equation 1

OpenStudy (anonymous):

what does the problem say to do with equation 2?

OpenStudy (anonymous):

same as we did in equation 1 but here we have to take negative values

OpenStudy (anonymous):

No, pay attention. The problem says transform equation 2 to its logarthmic form

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

im jst getting confused

OpenStudy (anonymous):

you just explain me

OpenStudy (anonymous):

i will follow it

OpenStudy (anonymous):

There are two parts of the problem. We finished the first part using equation 1. Now the second part of the problem says transform the second equation to logarithm form

OpenStudy (anonymous):

the second equation is X=5y^-2

OpenStudy (anonymous):

so you transform it to logarithms by taking log on both sides. that is log X = log(5y^-2)

OpenStudy (anonymous):

that is all I can explain. You should be able to take it from there.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

you are welcome.

OpenStudy (anonymous):

you helped me lot